OpenStudy (anonymous):

Write a polynomial function of minimum degree with real coefficients whose zeros include those listed. Write the polynomial in standard form. 4, -8, and 2 + 3i

OpenStudy (anonymous):

(x-4) (x+8) (x-(2+3i)) (x-(2-3i)) (x^2+4x-32) (4x^2+9) =4x^4+16x^3-119x^2+36x-288

OpenStudy (anonymous):

Remember factors including imaginary come in pairs so if you have 2i you must also have -2i, factor form it would be (x-2i)(x+2i)

OpenStudy (anonymous):

f(x) = x4 - 6x3 + 12x2 - 90x + 416 f(x) = x4 - 45x2 + 180x - 416 f(x) = x4 - 35x2 + 180x - 416 f(x) = x4 - 6x3 - 12x2 + 90x - 416 @ehuman

OpenStudy (anonymous):

those are your possible answers?

OpenStudy (anonymous):

Yes @ehuman

OpenStudy (anonymous):

I obviously made a mistake. Here are your 4 factors (x-4) (x+8) (x-(2+3i)) (x-(2-3i)) Foil them to find expanded solution.

OpenStudy (anonymous):

can you help me do that?

OpenStudy (anonymous):

@ehuman

OpenStudy (anonymous):

Apologies, I've worked it a couple different ways and don't come up with any of your choices. I must be rusty doing something wrong.

Aww):